### v1

@@ 3.4 Visualization Analysis of Experimental Results
Figure 2: Visualization of the model’s attention regions using saliency map overlaid on MRI slices. The red-highlighted areas indicate the regions most relevant to Alzheimer’s disease classification in axial, coronal, and sagittal views.
4 Conclusion
@@
We introduced the PGAD framework to address the challenges of aligning missing modality features and improving knowledge transfer in Alzheimer’s Disease diagnosis. By integrating Prototype Consistency Matching (PCM) and Adaptive Multi-Modal Sampling (AMS), PGAD effectively leverages incomplete multi-modal data, mitigating feature misalignment and enhancing cross-modal learning. Experimental results and ablation studies validate its effectiveness over existing methods.These findings highlight the potential of prototype-guided learning and adaptive sampling for improving multi-modal medical imaging under real-world missing data conditions.

### v2

@@ III-D2 Biomarker Identification via Grad-CAM
Figure 3: Grad-CAM visualization of the model’s attention regions on MRI slices. The red-highlighted areas indicate regions most relevant to AD classification in axial, coronal, and sagittal views.
IV Conclusion
@@
We introduced the PGAD framework to address the challenges of aligning missing modality features and improving knowledge transfer in Alzheimer’s Disease diagnosis. By integrating Prototype Consistency Matching (PCM) and Adaptive Multi-Modal Sampling (AMS), PGAD effectively leverages incomplete multi-modal data, mitigating feature misalignment and enhancing cross-modal learning. Experimental results demonstrate that PGAD achieves state-of-the-art performance in both AD classification and MCI conversion prediction under various missing rates. Ablation studies confirm the effectiveness of our proposed components. These findings highlight the potential of prototype-guided learning and adaptive sampling for improving multi-modal medical imaging under real-world missing data conditions. Future work will focus on extending the PGAD framework to incorporate three or more modalities and evaluating its performance in more complex, multi-center clinical scenarios.